Restart Night #10 UT 31 Aug 2020

OSA: S. Allanson

ISp: G. Bechetti

Mtn Manager: B. Bingham

SO: J. Power

Special Guests: D.Thompson on LUCI, S.Ertel on LMIRCam


Summary:

LBTI AO checkout completed. Pased with flying colours. We were able to perform a rough focus on LMIRcam with the ARGOS source as well. Dave took the data for the LUCI1 wavelength calibration. I verified the mirroring of mask 921459.VIPE1

Problems:

  1. LUCI-2 CCR offline - unavailable as of afternoon

Proposed Plan:

  • LUCI-1 - CRITICAL PRIORITY
  • PFS Alignment Alignment - daytime - into handover if needed - CRITICAL
  • Focus check time permitting (or if weather keeps us closed at sunset)
  • L! Rotator Center (0.5 hr) - high
    • M92, M34, or M52 - fields available for all night
  • L1 Rotator ZP for AGW - (1 hr) - critical
    • Long Slit Method - instructions to be provided by J. Power
      • Suggest doing 3 "masks" to verify the RotZP (based on experience with issues from L2 run)
  • AGW Transform for both LUCIs (use the N30 fieldstop, H or FeII filters) - (2 hr) - high/critical
  • LUCI-1 Imaging check (can run LUCI-2 in parallel to repeat) - high
    • Important to test in and out of field to "stress-test" AGW-1 - does it recover from long throws? and multi-point dithers?
    • Sidereal
      • Dither in-field
        • J1601 (sets around 10pm) - dithering in the field -Kband only
        • NGC 7027 (sets 4am) - cycles through all filters, can stop early
      • Dither out-of-field
        • M52 - (all night) - KJH
        • NGC 40 - (all night) - all filters
    • NS
      • Comet 29/P - (after 10pm) - JHK - need to update for tonight's ephemeris
  • Twilight Flats - low
  • Photometric Standards - (1-2hrs) - medium - only if clear - any seeing
  • AFTER LUCI-1 checkout:
  • LBCs
    • Collimate WRS/DOFPIA - high (4 x 0.25 hrs) - to check LBC-Blue
    • Superfocus sequence -PA=0,180 high (1 hr) - to check LBC-Blue
    • Rotator Center Checks, start with PA=0, 180 (track mode), then 90,270 - high (1hr) - to check LBC-Blue
    • Test Science OB - guiding - medium (1hr)
      • Sidereal
      • Non-Sidereal
    • Standard Photometric fields - requires clear skies - any seeing
    • Spectrophotometric Standards - requires clear skies - any seeing
    • Twlight Flats
  • MODS
    • Spectrophotometric stars if clear (and if not clear only if taking data)
    • Photometric Standards - (1 hr) - only if clear - any seeing
    • Mixed Mode LBC-R/MODS-2
    • Skyflats MODS2-Blue
    • Revisit Transient - (1hr) - Med/High
      • Same spectroscopy scripts (run 2x extra eposures - total of 4x900sec each channel)
      • If too faint for spectroscopy, take u,g,r,i,z imaging - script TBD

Closed Dome Activities

  • LUCI-Bino Flexure
  • DONELUCI1 Wavelength Calibration
  • DONELBTI AO Checkout
  • SOUL LUCI-1/AO Checkout

1:15 Conditions on the summit: Heavy cloud cover, drizzles earlier. Setting up to run LBTI WFS and AO prerun checkout closed dome

1:32 ARGOS light has been deployed, chamber lights off, in retro mode with M2 and M3 set to last position for LBTI WFS alignment.

  • Executing the
    LBTI Checkout: SX AdSec
    https://docs.google.com/spreadsheets/d/1NKQxa0I7EDg1spfKgGOTaOy69zN9MAKZkcMgddbpaWA/edit?usp=sharing
    Usage on /local/aomeas on lbtisxwfs6 is high at 84% and should be cleared out before run.
  • Aligning on M2 M3
    M2: 2.2983, 0.7501, -1.550, 245.0277, -209.0703, 0.
  • Loaded KL20 modal basis, usual reconstructions, and LBTI forgetting factor without issue.
    KL_20
    Rec_LBTI_IIR_b1_500m1000.fits
    Filtering: LBTI_IIR_b1_500m1000
    Closed loop bin1 500 modes frequency 1000 on argos source, optimized gains. Camera lens range good: 9,-21

    KL_20
    Rec_20180719_051224
    Filter forgettingFactorSOUL
    Closed loop bin 4, 54 modes, frequency 1000, on argos source. Camera lens range good -10, -45
  • Steve Ertel has LMIRCAM up and running to check image focus.
WFS disconnected from msgD, stopped and restarted the wfs arbitrator
  • Updating the Bayside stage positions to 52.8, 24.3 to center the argos source on the center of the LMIRCAM patrol field
    M2: 2.4642, 0.2025, -1.5430,390.32, -162.46 0
  • Ran focus sequence with ARGOS source and found optimal focus was around 77.5. Set the
    Bayside Z focus in the LBTI_ACE-AO in /home/sxwunit/aoroot/calib/wfs/current/setups to 78 as a starting point for an onsky focus sequence on a point source.
  • Closed with several disturbance files and reconstructors, no issues.
Dave is doing some LUCI1 wavelength work while we work on LBTI AO checkout. He gathered the LUCI2 wavelength data yesterday.

4:07 Steve is deploying the troughs preemptively before Bruce heads to bed.

4:33 Preset AO with AOS commands gui (manually set WFS Spec to LBTIWFS in Flao GUI). Intervention not set and closed and optimized without issue.

Tested a few setups, disturbance files, to be thorough but all looks just as happy as it did prior to shutdown.

For the books:

Baysides (xyz) 52.8,24.3,77.5 (align and focus on lmircam)
M2 Platform position Absolute (2.46,0.2,-1.53,386,-169,0)
M3 Absolute (-300,-25,0,4.2)

5:00 Done with the Adsec and LBTI wfs. Warmed the camera and powered off, w_stop. Adsec remains set. Also the rr_mode_enable was activated when I started tonight and I will leave it like that, unless a miracle happens and it looks like we will open.

Will park the ARGOS source once Dave hits a logical break in his scripts and I can turn the lights on (20 min or so).

7:20 Dave finished LUCI1 calibrations

8:50 Checking out some luci masks

  • luci1.20200831.0217 921459.VIPE1 in FPU. Confirmed mask is mirrored
  • luci1.20200831.0218 moving mask 931032.VIPB4 into the FPU slot 14 and another of the INAF masks. Verified no Xflip for this mask as expected it is just the VIPE1 mask.
  • put no mask in use and luci to bed.

-- BarryRothberg - 30 Aug 2020
  • Mask Snap of INAF mask luci1.8.VIPE.lms showing gerber mirrored:
    MaskSnap_921459.VIPE1.png
Topic attachments
I Attachment Action Size Date Who CommentSorted descending
MaskSnap_921459.VIPE1.pngpng MaskSnap_921459.VIPE1.png manage 2 MB 31 Aug 2020 - 09:13 JenniferPower Mask Snap of INAF mask luci1.8.VIPE.lms showing gerber mirrored
Topic revision: r4 - 31 Aug 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback